# Anharmonicity and Coulomb Pseudopotential Effects in YH6 and YH9 Superconductivity

### v1

@@ I Introduction
Since the discovery of superconductivity in mercury, efforts have been focused on searching for superconductors with higher superconducting transition temperatures ($T_{c}$). The Bardeen-Cooper-Schrieffer (BCS) theory explains conventional superconductivity through electron-phonon coupling (EPC). It also illuminates the path towards higher $T_{c}$, which requires stronger EPC and higher phonon frequency. In this regard, hydrides are excellent candidates for achieving high $T_{c}$ due to the light mass of hydrogen atoms. Guided by this idea, various hydride superconductors have been theoretically predicted [1]. Over the past decade, multiple hydride superconductors, such as H3[2, 3], LaH10 [4, 5, 6], CaH6 [7, 8, 9], YH6 and YH9 [10, 11, 12, 13], have been successfully synthesized and exhibit $T_{c}$ exceeding 200 K, while Li2MgH16 [14], which is predicted to possess the highest $T_{c}$, has yet to be confirmed in experiments.
For hydride superconductors, discrepancies often arise between experimental and theoretical $T_{c}$ values. Several explanations have been proposed, including anharmonic effects on ion motion [15, 16, 11], diffusion of ion vacancies [17, 18], structure phase transitions [19, 20, 21], and nonhydrostatic stress in experiments [22]. Among these mechanisms, the quantum and anharmonic effects of hydrogen ions are widely accepted as the cause for the overestimation of $T_{c}$ in theoretical calculations. One well-established method to incorporate these effects is the stochastic self-consistent harmonic approximation (SSCHA) [23, 24, 25], which has been highly successful in predicting $T_{c}$’s for various hydride superconductors [15, 16].
@@
In this paper, we present a first-principles study of anharmonic effects in yttrium hydride superconductors. To account for the anharmonicity and quantum fluctuations of hydrogen ions, we adopt the stochastic path-integral approach (SPIA) implemented with density functional theory (DFT) [27, 28, 29]. SPIA is a non-perturbative approach that does not make assumptions about the nature of ion motion or the existence of an effective harmonic potential. Compared to SSCHA, SPIA has broader applicability. It has been successfully applied to metallic hydrogen liquid [27], anharmonic solid H3[29], superionic superconductor Li2MgH16 [30], and hydrogen-deficient LaH10-δ [17]. Using SPIA, we investigate the $T_{c}$ of YH6 at 165 GPa and YH9 at 255 GPa. We find that anharmonic effects significantly alter EPC parameters and increase the average phonon frequency, aligned with prior SSCHA findings. By properly taking into account the renormalization of the Coulomb pseudopotential due to frequency cutoff, which is neglected in earlier studies, we obtain $T_{c}$ values that closely match experimental observations for both compounds. This indicates that anharmonic effects are sufficient to explain the discrepancies between experimental and theoretical results in these systems.
@@ II Stochastic path-integral approach
SPIA is a method for determining $T_{c}$ of superconductors based on ab initio path integral molecular dynamics (PIMD) [31, 32]. The key to this approach is to determine the effective interaction between electrons through the fluctuation of the electron-ion scattering $T$ matrix.
@@
The propagation and pairing of electrons occur in states that diagonalize the physical electron Green’s function. For harmonic solids, this basis is just the Bloch states. For anharmonic solids, a generalized Bloch basis can be determined by diagonalizing the effective Hamiltonian which is derived from the electron Green’s function in plane-wave basis [29]. The index of the generalized Bloch states can be written as $1=(n,\boldsymbol{k},\omega_{j})$, where $\boldsymbol{k}$ is the quasi-wave-vector, $n$ is the band index, and $\omega_{j}$ is the Fermionic Matsubara frequency.
@@
For conventional superconductors, it is usually reasonable to apply the isotropic approximation [33]. Under this approximation, we replace the effective interaction $\hat{W}$ with an isotropic average over the Fermi surface, which is known as the EPC parameter $\lambda$ given by
@@
where $N(\epsilon_{F})$ is the electron density of states (DOS) at the Fermi level, and $\langle\cdots\rangle_{\mathrm{FS}}$ denotes the average for all initial states $(n\boldsymbol{k})$ and final states $(n^{\prime}\boldsymbol{k}^{\prime})$ over the Fermi surface. The EPC parameters then enter the isotropic linearized Eliashberg equation [34] which we use to determine $T_{c}$ in this paper
@@
where $\tilde{\omega}(j)$ is the renormalized Matsubara frequency under isotropic approximation given by $\tilde{\omega}(j)=\frac{\pi}{\hbar\beta}(2j+1+\lambda(0)+2\sum_{i=1}^{j}\lambda(i))$. To include the effect of the Coulomb interaction between electrons, we introduce the Morel-Anderson pseudopotential $\mu^{*}$ [35]. Eq. (3) is an eigenvalue problem to be solved for the eigenvalue $\rho$ and renormalized gap $\Delta_{j}$. The emergence of a positive $\rho$ indicates the instability towards forming Cooper pairs and thus the superconducting phase. It is shown in Ref. [27] that the linearized Eliashberg equation remains valid for general systems including anharmonic solids as long as the effective electron-electron interaction is determined by Eq. (1) when calculating the EPC parameters.
When evaluating the EPC parameters in Eq. (2), $N(\epsilon_{F})$ is determined by averaging the DOS over all PIMD configurations on a denser $\boldsymbol{k}$-grid to obtain converged values (see Appendix A). The average of $\hat{W}$ is weighted by the Lorentzian function $\gamma/[(\epsilon_{n\boldsymbol{k}}-\epsilon_{F})^{2}+\gamma^{2}]$, where $\epsilon_{n\boldsymbol{k}}$ is the eigen-energy of the generalized Bloch states. The half-width of the Lorentzian is set to $\gamma=\tilde{\omega}(0)=\frac{\pi}{\hbar\beta}(1+\lambda(0))$ [27].
When solving the linearized Eliashberg equation Eq. (3), we need to set a truncation $N$ for the summation over $j$. This truncation $N$ (i.e., the frequency cutoff) results in a renormalization of $\mu^{*}$ [34] given by
$\frac{1}{\mu^{*}(N)}=\frac{1}{\mu^{*}}+\ln\frac{\bar{\omega}_{2}}{\omega_{N}},$ (4)
where the average phonon frequency $\bar{\omega}_{2}$ is determined by the asymptotic behavior of $\lambda(m)$ through $\bar{\omega}_{2}=\lim_{m\to\infty}\frac{2\pi}{\hbar\beta}\sqrt{\frac{m^{2}\lambda(m)}{\lambda(0)}}$. This renormalized $\mu^{*}(N)$, which replaces $\mu^{*}$ in Eq. (3), is crucial for obtaining a $T_{c}$ that is independent of the truncation $N$ (see Sec. III.3).
@@ III.1 Computational details
To accelerate the PIMD simulation while maintaining its first-principles accuracy, we apply the machine learning force field (MLFF) module of VASP [39, 40, 41]. The modified program allows an on-the-fly training of MLFF in (PI)MD simulations [30]. The details of training can be found in Appendix B.
Using the MLFF generated above, we perform PIMD simulations for YH6 and YH9 with bead number $N_{b}=16$. A Langevin thermostat with the friction coefficients of centroid mode set to [10 10] ps-1 is used to control the temperature of the canonical (NVT) ensemble [42]. We use a $3\times 3\times 4$ supercell containing 252 atoms for YH6 and a $3\times 3\times 2$ supercell containing 360 atoms for YH9. YH6 is simulated at 220 K while YH9 is simulated at 240 K. The overall simulation time is 3 ps with a time step of 0.5 fs.
The ion configurations of YH6 and YH9 are uniformly sampled with a spacing of 40 time steps after a 0.5-ps equilibrium. DFT calculations are performed for these configurations. An energy cutoff of 350 eV for plane waves is used to expand electron wave functions, and a $3\times 3\times 3$ $\Gamma$-centered $\boldsymbol{k}$-point grid is used to sample the Brillouin zone of the supercell. The converged DFT results are then used as inputs of our MATLAB implementation of SPIA [29] to reconstruct the Hamiltonian. Finally, the effective electron-electron interaction $\hat{W}$ is calculated on the irreducible wedge of a $6\times 6\times 6$ $\boldsymbol{k}$-point grid of the supercell. The EPC parameters $\lambda(m)$ are calculated by averaging $\hat{W}$ near the Fermi surface. The half-width of the Lorentzian smearing is 0.16 eV for YH6 and 0.18 eV for YH9.
@@ III.2 SPIA results
The EPC parameters of YH6 and YH9 are shown in Fig. 2. For YH6, the SPIA results exhibit a significant suppression of $\lambda(0)$, $\lambda(1)$ and $\lambda(2)$ compared to the DFPT results. The values of these parameters, the behavior of $\bar{\omega}_{2}(m)=2\pi/\hbar\beta\sqrt{m^{2}\lambda(m)/\lambda(0)}$ and its asymptotic value (i.e., the average phonon frequency $\bar{\omega}_{2}$) are close to the results of SSCHA, which is shown in Fig. 2(a) and Table 1.
Figure 2: (a) EPC parameters $\lambda(m)$ of YH6 under 165 GPa at 220 K. DFPT and SSCHA results are calculated using the Eliashberg spectral functions presented in Ref. [11]. (b) EPC parameters $\lambda(m)$ of YH9 under 255 GPa at 240 K. DFPT result is calculated using the Eliashberg spectral function presented in Ref. [20] under 260 GPa. The inset shows the asymptotic behavior of $\bar{\omega}_{2}(m)=2\pi/\hbar\beta\sqrt{m^{2}\lambda(m)/\lambda(0)}$, whose asymptotic value gives the average phonon frequency $\bar{\omega}_{2}$.
System Method $\lambda(0)$ $\lambda(1)$ $\lambda(2)$ $\bar{\omega}_{2}$ (K) $T_{c}$ (K)
@@
SPIA 1.749 0.944 0.460 1905 238
Table 1: The first few EPC parameters $\lambda(m)$, average phonon frequency $\bar{\omega}_{2}$ and predicted $T_{c}$ of YH6 and YH9. We set $\mu^{*}=0.11$ for all methods. For DFPT and SSCHA results, their EPC parameters and average phonon frequencies are calculated based on the Eliashberg spectral functions presented in Refs. [11, 20]. The $T_{c}$ values are also directly taken from the same sources. $T_{c}$ values in parentheses are calculated using our method with renormalized $\mu^{*}(N)$. The DFPT result for YH9 is for 260 GPa.
For YH9, the SPIA results exhibit a strong suppression of $\lambda(0)$, a slight suppression of $\lambda(1)$ and an enhancement of $\lambda(m)\ (m\geq 2)$ compared to the DFPT results. This leads to a notable enhancement of $\bar{\omega}_{2}(m)$ values and thus the average phonon frequency $\bar{\omega}_{2}$, as shown in Fig. 2(b) and Table 1. While the suppressed $\lambda(1)$ would lower $T_{c}$, its effect is compensated by the enhancement of $\lambda(m)\ (m\geq 2)$. This could explain why DFPT still yields a decent $T_{c}$ despite the sizable anharmonic corrections.
Setting $\mu^{*}=0.11$ [33], we calculate $T_{c}$ as 209 K for YH6 and 238 K for YH9. Table 1 lists the results obtained using different methods. Figure 3 summarizes observed or predicted $T_{c}$ values from various experiments and calculations. It can be seen that our SPIA predictions agree well with the experimentally observed $T_{c}$ values.
Figure 3: $T_{c}$ values of YH6 and YH9 at various pressures. Filled points indicate experimental data from Refs. [11, 12, 13]. Empty points represent theoretical predictions from Refs. [11, 20] and our SPIA calculations. $\mu^{*}=0.11$ is used in all calculations.
@@ III.3 Renormalization of the Coulomb pseudopotential
Properly taking into account the renormalization of $\mu^{*}$ is crucial for correctly evaluating $T_{c}$. The renormalization of $\mu^{*}$ due to the frequency cutoff in the Eliashberg equation has been discussed in Sec. II. In Fig. 4, we test the relation between $T_{c}$ and frequency cutoff by solving the linearized Eliashberg equation with and without the $\mu^{*}$ renormalization using EPC parameters of YH6 calculated by SSCHA in Ref. [11]. It can be seen that without the $\mu^{*}$ renormalization, $T_{c}$ increases continuously with the frequency cutoff and does not show a trend of convergence even at large $N$ values. In contrast, with the $\mu^{*}$ renormalization, $T_{c}$ converges quickly. This indicates that only with the renormalization can we obtain a $T_{c}$ that is independent of the truncation.
Figure 4: The relation between $T_{c}$ and frequency cutoff when solving the linearized Eliashberg equation. The truncation $N$ denotes the frequency cutoff at $\omega_{N}=(2N+1)\pi/\hbar\beta$. The blue line is calculated with the bare $\mu^{*}=0.11$ while the red line is calculated with the renormalized $\mu^{*}(N)$ defined in Eq. (4). The Eliashberg spectral function used in calculation is the SSCHA result of YH6 from Ref. [11].
The renormalization of $\mu^{*}$ plays a decisive role in our accurate predictions of $T_{c}$’s. As shown in Table 1, SPIA yields EPC parameters that are close to the SSCHA results for YH6, yet it leads to a significant suppression of $T_{c}$. This can be attributed to the renormalization, which increases $\mu^{*}$. Table 1 also includes $T_{c}$’s calculated using the renormalized $\mu^{*}$ for both DFPT and SSCHA methods (values in parentheses). These results reveal that, had the renormalization of $\mu^{*}$ been considered, SSCHA would have made a satisfactory prediction for the $T_{c}$ of YH6. This suggests that anharmonic effects, combined with $\mu^{*}$ renormalization, are sufficient to bridge the gap between theory and experimental observations.
IV Summary

### v2

# Anharmonicity and Coulomb pseudopotential effects on superconductivity in YH6 and YH9

@@ I Introduction
Since the discovery of superconductivity in mercury, efforts have been focused on searching for superconductors with higher superconducting transition temperatures ($T_{c}$). The Bardeen-Cooper-Schrieffer (BCS) theory explains conventional superconductivity through electron-phonon coupling (EPC). It also illuminates the path towards higher $T_{c}$, which requires stronger EPC and higher phonon frequency. In this regard, hydrides are excellent candidates for achieving high $T_{c}$ due to the light mass of hydrogen atoms. Guided by this idea, various hydride superconductors have been theoretically predicted [hydride_review]. Over the past decade, multiple hydride superconductors, such as H3[H3S_the, H3S_exp], LaH10 [LaH10_the, LaH10_exp1, LaH10_exp2], CaH6 [CaH6_the, CaH6_exp1, CaH6_exp2], YH6 and YH9 [YH_the1, YH6_exp1, YH_exp2, YH_exp3], have been successfully synthesized and exhibit $T_{c}$ exceeding 200 K, while Li2MgH16 [LMH_the], which is predicted to possess the highest $T_{c}$, has yet to be confirmed in experiments.
For hydride superconductors, discrepancies often arise between experimental and theoretical $T_{c}$ values. Several explanations have been proposed, including anharmonic effects on ion motion [H3S_SSCHA, CaH6_SSCHA, YH6_exp1], diffusion of ion vacancies [LaH10_SPIA, CaH6_SPIA], structure phase transitions [CaH6_phase, YH9_phase, LaH10_phase], and nonhydrostatic stress in experiments [YH6_stress]. Among these mechanisms, the quantum and anharmonic effects of hydrogen ions are widely accepted as the cause for the overestimation of $T_{c}$ in theoretical calculations. One well-established method to incorporate these effects is the stochastic self-consistent harmonic approximation (SSCHA) [SSCHA_1, SSCHA_2, SSCHA_3], which has been highly successful in predicting $T_{c}$’s for various hydride superconductors [H3S_SSCHA, CaH6_SSCHA].
@@
In this paper, we present a first-principles study of anharmonic effects in yttrium hydride superconductors. To account for the anharmonicity and quantum fluctuations of hydrogen ions, we adopt the stochastic path-integral approach (SPIA) implemented with density functional theory (DFT) [SPIA, H_SPIA, H3S_SPIA]. SPIA is a non-perturbative approach that does not make assumptions about the nature of ion motion or the existence of an effective harmonic potential. Compared to SSCHA, SPIA offers broader applicability. It can be applied not only to solid systems [H3S_SPIA], but also to systems with diffusive atoms, such as liquids [SPIA], superionic solids [LMH_SPIA], and systems with diffusive defects [LaH10_SPIA]. Using SPIA, we investigate the $T_{c}$ of YH6 at 165 GPa and YH9 at 255 GPa. We find that anharmonic effects significantly alter EPC parameters and increase the average phonon frequency, aligned with prior SSCHA findings. By properly taking into account the renormalization of the Coulomb pseudopotential due to frequency cutoff, which is neglected in earlier studies, we obtain $T_{c}$ values that closely match experimental observations for both compounds. This indicates that anharmonic effects are sufficient to explain the discrepancies between experimental and theoretical results in these systems.
@@ II Stochastic path-integral approach
SPIA is a method for determining $T_{c}$ of superconductors based on ab initio path integral molecular dynamics (PIMD) [PIMD, AI_PIMD]. The key to this approach is to determine the effective interaction between electrons through the fluctuation of the electron-ion scattering $T$ matrix.
@@
The propagation and pairing of electrons occur in states that diagonalize the physical electron Green’s function. For harmonic solids, this basis is just the Bloch states. For anharmonic solids, a generalized Bloch basis can be determined by diagonalizing the effective Hamiltonian which is derived from the electron Green’s function in plane-wave basis [H3S_SPIA]. The index of the generalized Bloch states can be written as $1=(n,\boldsymbol{k},\omega_{j})$, where $\boldsymbol{k}$ is the quasi-wave-vector, $n$ is the band index, and $\omega_{j}$ is the Fermionic Matsubara frequency.
@@
For conventional superconductors, it is usually reasonable to apply the isotropic approximation [YH_the2]. Under this approximation, we replace the effective interaction $\hat{W}$ with an isotropic average over the Fermi surface, which is known as the EPC parameter $\lambda$ given by
@@
where $N(\epsilon_{F})$ is the electron density of states (DOS) at the Fermi level, and $\langle\cdots\rangle_{\mathrm{FS}}$ denotes the average for all initial states $(n\boldsymbol{k})$ and final states $(n^{\prime}\boldsymbol{k}^{\prime})$ over the Fermi surface. The EPC parameters then enter the isotropic linearized Eliashberg equation [Allen-Dynes] which we use to determine $T_{c}$ in this paper
@@
where $\tilde{\omega}(j)$ is the renormalized Matsubara frequency under isotropic approximation given by $\tilde{\omega}(j)=\frac{\pi}{\hbar\beta}(2j+1+\lambda(0)+2\sum_{i=1}^{j}\lambda(i))$. To include the effect of the Coulomb interaction between electrons, we introduce the Morel-Anderson pseudopotential $\mu^{*}$ [mustar, mustar_GW]. Eq. (3) is an eigenvalue problem to be solved for the eigenvalue $\rho$ and renormalized gap $\Delta_{j}$. The emergence of a positive $\rho$ indicates the instability towards forming Cooper pairs and thus the superconducting phase. It is shown in Ref. [SPIA] that the linearized Eliashberg equation remains valid for general systems including anharmonic solids as long as the effective electron-electron interaction is determined by Eq. (1) when calculating the EPC parameters.
When evaluating the EPC parameters in Eq. (2), $N(\epsilon_{F})$ is determined by averaging the DOS over all PIMD configurations on a denser $\boldsymbol{k}$-grid to obtain converged values (see Appendix A). The average of $\hat{W}$ is weighted by the Lorentzian function $\gamma/[(\epsilon_{n\boldsymbol{k}}-\epsilon_{F})^{2}+\gamma^{2}]$, where $\epsilon_{n\boldsymbol{k}}$ is the eigen-energy of the generalized Bloch states. The half-width of the Lorentzian is set to $\gamma=\tilde{\omega}(0)=\frac{\pi}{\hbar\beta}(1+\lambda(0))$ [SPIA].
When solving the linearized Eliashberg equation Eq. (3), we need to set a truncation $N$ for the summation over $j$. This truncation $N$ (i.e., the Matsubara frequency cutoff) results in a renormalization of $\mu^{*}$, as pointed out by Allen and Dynes [Allen-Dynes]. It is given by
$\frac{1}{\mu^{*}(N)}=\frac{1}{\mu^{*}}+\ln\frac{\nu_{\mathrm{ph}}}{\omega_{N}},$ (4)
where $\nu_{\mathrm{ph}}$ is the phonon cutoff frequency at which $\mu^{*}$ is evaluated [McMillan, mustar_GW]. This renormalized $\mu^{*}(N)$, which replaces $\mu^{*}$ in Eq. (3), is crucial for obtaining a $T_{c}$ that is independent of the truncation $N$ (see Sec. III.3).
@@ III.1 Computational details
To accelerate the PIMD simulation while maintaining its first-principles accuracy, we apply the machine learning force field (MLFF) module of VASP [MLFF_1, MLFF_2, MLFF_3]. The modified program allows an on-the-fly training of MLFF in (PI)MD simulations [LMH_SPIA]. The details of training can be found in Appendix B.
Using the MLFF generated above, we perform PIMD simulations for YH6 and YH9 with bead number $N_{b}=16$. A Langevin thermostat with the friction coefficients of centroid mode set to [10 10] ps-1 is used to control the temperature of the canonical (NVT) ensemble [Langevin]. We use a $3\times 3\times 4$ supercell containing 252 atoms for YH6 and a $3\times 3\times 2$ supercell containing 360 atoms for YH9. YH6 is simulated at 220 K while YH9 is simulated at 240 K. The overall simulation time is 3 ps with a time step of 0.5 fs.
The ion configurations of YH6 and YH9 are uniformly sampled with a spacing of 40 time steps after a 0.5-ps equilibrium. DFT calculations are performed for these configurations. An energy cutoff of 350 eV for plane waves is used to expand electron wave functions, and a $3\times 3\times 3$ $\Gamma$-centered $\boldsymbol{k}$-point grid is used to sample the Brillouin zone of the supercell. The converged DFT results are then used as inputs of our MATLAB implementation of SPIA [H3S_SPIA] to reconstruct the Hamiltonian. Finally, the effective electron-electron interaction $\hat{W}$ is calculated on the irreducible wedge of a $6\times 6\times 6$ $\boldsymbol{k}$-point grid of the supercell. The EPC parameters $\lambda(m)$ are calculated by averaging $\hat{W}$ near the Fermi surface. The half-width of the Lorentzian smearing is 0.16 eV for YH6 and 0.18 eV for YH9.
@@ III.2 SPIA results
The EPC parameters of YH6 and YH9 are shown in Fig. 2. For YH6, the SPIA results exhibit a significant suppression of $\lambda(0)$, $\lambda(1)$ and $\lambda(2)$ compared to the DFPT results. The values of these parameters, the behavior of $\bar{\nu}_{2}(m)=2\pi/\hbar\beta\sqrt{m^{2}\lambda(m)/\lambda(0)}$ and its asymptotic value (i.e., the average phonon frequency $\bar{\nu}_{2}$) are close to the results of SSCHA, which is shown in Fig. 2(a) and Table 1.
Figure 2: (a) EPC parameters $\lambda(m)$ of YH6 under 165 GPa at 220 K. DFPT and SSCHA results are calculated using the Eliashberg spectral functions presented in Ref. [YH6_exp1]. (b) EPC parameters $\lambda(m)$ of YH9 under 255 GPa at 240 K. DFPT result is calculated using the Eliashberg spectral function presented in Ref. [YH9_phase] under 260 GPa. The inset shows the asymptotic behavior of $\bar{\nu}_{2}(m)=2\pi/\hbar\beta\sqrt{m^{2}\lambda(m)/\lambda(0)}$, whose asymptotic value gives the average phonon frequency $\bar{\nu}_{2}$.
System Method $\lambda(0)$ $\lambda(1)$ $\lambda(2)$ $\bar{\nu}_{2}$ (K) $T_{c}$ (K)
@@
SPIA 1.749 0.944 0.460 1905 240
Table 1: The first few EPC parameters $\lambda(m)$, average phonon frequency $\bar{\nu}_{2}$ and predicted $T_{c}$ of YH6 and YH9. We set $\mu^{*}=0.11$ for all methods. For DFPT and SSCHA results, their EPC parameters and average phonon frequencies are calculated based on the Eliashberg spectral functions presented in Refs. [YH6_exp1, YH9_phase]. The $T_{c}$ values are also directly taken from the same sources. $T_{c}$ values in parentheses are calculated using our method with renormalized $\mu^{*}(N)$. The DFPT result for YH9 is for 260 GPa.
For YH9, the SPIA results exhibit a strong suppression of $\lambda(0)$, a slight suppression of $\lambda(1)$ and an enhancement of $\lambda(m)\ (m\geq 2)$ compared to the DFPT results. This leads to a notable enhancement of $\bar{\nu}_{2}(m)$ values and thus the average phonon frequency $\bar{\nu}_{2}$, as shown in Fig. 2(b) and Table 1. While the suppressed $\lambda(1)$ would lower $T_{c}$, its effect is compensated by the enhancement of $\lambda(m)\ (m\geq 2)$. This could explain why DFPT still yields a decent prediction of $T_{c}$ despite the sizable anharmonic corrections.
Figure 3: $T_{c}$ values of YH6 and YH9 at various pressures. Filled points indicate experimental data from Refs. [YH6_exp1, YH_exp2, YH_exp3]. Empty points represent theoretical predictions from Refs. [YH6_exp1, YH9_phase] and our SPIA calculations. $\mu^{*}=0.11$ is used in all calculations.
Setting $\mu^{*}=0.11$, as determined from $GW$ methods [YH_the2, mustar_GW], and $\nu_{\mathrm{ph}}=2518$ K [mustar_GW], we find $T_{c}$ to be 212 K for YH6 and 240 K for YH9. Table 1 lists the results obtained using different methods. Figure 3 summarizes observed or predicted $T_{c}$ values from various experiments and calculations. It can be seen that our SPIA predictions agree well with the experimentally observed $T_{c}$ values.
@@ III.3 Renormalization of the Coulomb pseudopotential
Properly taking into account the renormalization of $\mu^{*}$ due to the frequency cutoff in the Eliashberg equation, as specified in Eq. (4), is crucial for correctly evaluating $T_{c}$ [Allen-Dynes]. In Fig. 4, we test the relation between $T_{c}$ and frequency cutoff by solving the linearized Eliashberg equation with and without the $\mu^{*}$ renormalization using EPC parameters of YH6 calculated by SSCHA in Ref. [YH6_exp1]. It can be seen that without the $\mu^{*}$ renormalization, $T_{c}$ increases continuously with the frequency cutoff and does not show a trend of convergence even at large $N$ values. In contrast, with the $\mu^{*}$ renormalization, $T_{c}$ converges quickly. This indicates that only with the renormalization can we obtain a $T_{c}$ that is independent of the truncation.
Figure 4: The relation between $T_{c}$ and frequency cutoff when solving the linearized Eliashberg equation. The truncation NN denotes the frequency cutoff at $\omega_{N}=(2N+1)\pi/\hbar\beta$. The blue line is calculated with the bare $\mu^{*}=0.11$ while the red line is calculated with the renormalized $\mu^{*}(N)$ defined in Eq. (4). The Eliashberg spectral function used in calculation is the SSCHA result of YH6 from Ref. [YH6_exp1].
The renormalization of $\mu^{*}$ plays an important role in our accurate predictions of $T_{c}$’s. As shown in Table 1, SPIA yields EPC parameters that are close to the SSCHA results for YH6, yet it leads to a significant suppression of $T_{c}$. This can be attributed to the renormalization, which increases $\mu^{*}$. Table 1 also includes $T_{c}$’s calculated using the renormalized $\mu^{*}$ for both DFPT and SSCHA methods (values in parentheses). These results reveal that, had the renormalization of $\mu^{*}$ been considered, SSCHA would have made a satisfactory prediction for the $T_{c}$ of YH6. This suggests that anharmonic effects, combined with $\mu^{*}$ renormalization, are sufficient to bridge the gap between theory and experimental observations.
IV Summary and discussions